Happyfeet180:

Lily is a botanist who works for a garden that many tourists visit. The function f(s) = 2s + 30 represents the number of flowers that bloomed, where s is the number of seeds she planted. The function s(w) = 40w represents the number of seeds she plants per week, where w represents the number of weeks. Part A: Write a composite function that represents how many flowers Lily can expect to bloom over a certain number of weeks. (4 points) Part B: What are the units of measurement for the composite function in Part A? (2 points) Part C: Evaluate the composite function in Part A for 35 weeks. (4 points)

Vocaloid:

part A) first function f(s) gives you flowers (f) per (s) seeds second function s(w) gives you seeds (s) per week (w) so if you want flowers (f) per week (w), you'll need to take the second function s, plug it into the first function, and simplify f(s) = 2s + 30, and since s(w) = 40w, replace "s" with "40w" in the first function

Vocaloid:

part B) since the function from part A) gives you the # of flowers for w weeks, the unit is simply # of flowers part C) plug in w = 35 into the function

Happyfeet180:

So, f(s(w))

Vocaloid:

yeah, that's the logic, just make sure to write out the whole function in terms of w

Happyfeet180:

f(s(w))=2(40w)+30 ??

Happyfeet180:

@vocaloid f(s(w))=2(40w)+30 ??

Happyfeet180:

@vocaloid or f(s(w))=80w+30

Vocaloid:

Yes, 80w + 30 is correct

Happyfeet180:

@Volcaloid for part c do i plug it in to f(s(w))=2(40w)+30 or f(s(w))=80w+30

Vocaloid:

both the functions you've written are equal. doesn't matter. 80w + 30 is easier to do.

Happyfeet180:

ok thank you